y''+y=sin^2(x)


1
Expert's answer
2021-10-03T17:43:52-0400

"y^{\prime\prime}+y=\sin^{2}(x)" (1)

The general solution will be the sum of the complementary solution and particular solution.

Find the complementary solution for the homogeneous equation:

"y^{\prime\prime}+y=0".

Compose the characteristic equation:

"\lambda^{2}+1=0\implies \lambda=i" or "\lambda=-i".

For complementary solution:

"y_{c}=y_{1}(x)+y_{2}(x)", where "y_{1}(x)=C_{1}\cos(x)" and "y_{2}(x)=C_{2}\sin(x)."

"y_{c}=C_{1}\cos(x)+C_{2}\sin(x)"

-------------------------------------------------------------------------------------------------------------------------------

Find the particular solution for (1).

"\sin^{2}(x)=\frac{1}{2}-\frac{1}{2}\cos(2x)\implies y^{\prime\prime}+y=\frac{1}{2}-\frac{1}{2}\cos(2x)".

It's clear that "y_{p}=\frac{1}{2}+\alpha\cos(2x)", "\alpha - constant".

Find "\alpha."

"y_p^{\prime\prime}+y_p=4\alpha(-\cos(2x))+\frac{1}{2}+\alpha\cos(2x)=\frac{1}{2}-\frac{1}{2}\cos(2x)".

"\boxed{\alpha=\frac{1}{6}}"

"y_{p}=\frac{1}{2}+\frac{1}{6}\cos(2x)".

The general solution is:

"y(x)=y_{c}(x)+y_{p}(x)=C_{1}\cos(x)+C_{2}\sin(x)+\frac{1}{2}+\frac{1}{6}\cos(2x)".

"\boxed{y=C_{1}\cos(x)+C_{2}\sin(x)+\frac{1}{2}+\frac{1}{6}\cos(2x)}"